# GroundingBooth: Grounding Text-to-Image Customization

### v1

@@ Abstract
Recent studies in text-to-image customization show great success in generating personalized object variants given several images of a subject. While existing methods focus more on preserving the identity of the subject, they often fall short of controlling the spatial relationship between objects. In this work, we introduce GroundingBooth, a framework that achieves zero-shot instance-level spatial grounding on both foreground subjects and background objects in the text-to-image customization task. Our proposed text-image grounding module and masked cross-attention layer allow us to generate personalized images with both accurate layout alignment and identity preservation while maintaining text-image coherence. With such layout control, our model inherently enables the customization of multiple subjects at once. Our model is evaluated on both layout-guided image synthesis and reference-based customization tasks, showing strong results compared to existing methods. Our work is the first work to achieve a joint grounding of both subject-driven foreground generation and text-driven background generation. The project page is available at https://groundingbooth.github.io.
Figure 1: We propose GroundingBooth, a framework for grounded text-to-image customization. GroundingBooth supports: (a) grounded single-subject customization, and (b) multi-subjects and text entities joint grounded customization, achieving joint grounding of subject-driven foreground and text-driven background generation with identity preservation and text-image alignment.
@@ 1 Introduction
Text-to-image customization, also known as subject-driven image generation or personalized text-to-image synthesis, is a task that requires a model to generate diverse variants of a subject given a set of images of the target subject. Text-to-image customization has achieved significant progress during the past few years, allowing for more advanced image manipulation. For example, the test-time-finetuning based methods like Dreambooth (Ruiz et al., 2023), Textual Inversion Gal et al. (2022), and Custom Diffusion Kumari et al. (2023) use a few images of the same object to finetune a pretrained diffusion model and generate variants of the object from input prompts. The encoder-based methods such as ELITE Wei et al. (2023) and InstantBooth Shi et al. (2023) eliminate test-time-finetuning by learning a generalizable image encoder and attention modules. Despite their success, existing personalization methods mainly focus on generating identity-preserved images from the input prompt and fail to accurately describe the spatial relationship of objects and backgrounds. In real-world scenarios of image customization, it is a crucial user need to achieve fine-grained and accurate layout control on each of the generated objects.
@@
To tackle this issue, we propose to deal with a more challenging task, grounded text-to-image customization, which extends the existing text-to-image customization task by enabling grounding controllability over both the foreground subjects and background objects. Specifically, under this new setting, the inputs usually include a prompt, images of subjects, and optional bounding boxes of the subjects and text entities. The model is required to generate text-aligned background objects and identity-preserved foreground subjects, while the spatial location of all the grounded objects and subjects are exactly aligned with the input bounding boxes. It is non-trivial to achieve all these effects simultaneously, as we are indeed handling multiple tasks together and it is challenging for the model to be generalizable to all sub-tasks.
@@
There are a few related works to our new task, which, however, show significant limitations. On the one hand, although existing grounded text-to-image diffusion models such as LayoutDiffusion (Zheng et al., 2023b) and GLIGEN Li et al. (2023) have made attempts at spatial controllability, they cannot achieve identity preservation of the subjects. On the other hand, subject-driven personalization methods mainly focus on the identity preservation of the reference objects, while limited attempts have been made on layout control of either subjects or background objects. There is another line of related works (Chen et al., 2023b; Song et al., 2022; 2024) that achieves customized image composition. They can control the location of the input subject under the image composition setting but are neither able to achieve text-to-image synthesis nor control the spatial location of the background objects.
@@
To fully address our new task, we propose GroundingBooth, a general framework for grounded text-to-image customization. Specifically, based on a pretrained text-to-image model, we build a new text-image joint grounding module that encourages both the foreground subjects and background objects to accurately follow the locations indicated by the input bounding boxes. To further enhance the identity preservation of the subjects, we propose a masked cross-attention layer in the U-Net, which disentangles subject-driven foreground generation and text-driven background generation in each transformer block, effectively preventing the blending of multiple objects in the same location and enforcing the generation of clear subjects. As shown in Fig. 1, with such dedicated designs of the network structures, our framework not only achieves grounded text-to-image customization with a single subject (Fig. 1 (a)), but also supports multi-subject customization (Fig. 1 (b)), where users can input multiple subjects along with their bounding boxes, and our model can generate each subject in the exact target region with identity preservation and scene harmonization. Our model also allows for the grounding of background objects (Fig. 1 (b)).
@@
  • We propose a general framework, GroundingBooth, that achieves grounded text-to-image customization. Specifically, our model achieves joint layout grounding of both the foreground subjects and the text-guided background, while maintaining accurate identity of the subjects. Furthermore, our model enables the customization of multiple subjects at once.
@@ 2.1 Subject-driven Image Generation
Subject-driven image generation, also known as personalized text-to-image generation or image customization, aims to generate target images based on customized objects and a text prompt that describes the target context  (Chen et al., 2023a; Pan et al., 2024; Xiao et al., 2023; Wang et al., 2024a). In this task, the specific identity of the input reference images is defined as a subject or a concept. So far, existing image customization works can be categorized into three major types. The first type is test-time-finetuning methods (Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023). These methods first finetune a pretrained diffusion model on a few subject images so that the model is adapted to a new identifier token representing the new concept. Then they generate new images from prompts containing the identifier. Such test-time finetuning is computationally intensive. The second type of method is encoder-based customization methods (Arar et al., 2023; Wei et al., 2023; Shi et al., 2023; Zhang et al., 2024), which eliminates the test-time finetuning by learning a generalizable diffusion model that can adapt to new subjects on a large-scale training data. The generalizable model usually contains image encoders that map the subject images into dense tokens and attention modules that integrate vision tokens with text tokens. These methods can achieve much faster image customization during inference, but they rely heavily on large-scale multi-view training data, and identity preservation may not be perfect in out-of-distribution cases. The third type of methods (Roich et al., 2021; Gal et al., 2023) is a combination of the first two methods, which first learn an image encoder to extract identity tokens of the input subject and then finetune the model on the subject images for a few steps.
Note that most existing subject-driven image generation methods are focusing on synthesizing personalized image variants from prompts. They show quite limited performance in controlling the layout of the generated scenes and modeling the spatial relationship between objects. On the contrary, our work performs well on not only generating identity-preserved, text-aligned images but also controlling the layout of both the subjects and background. A recent related work Anydoor Chen et al. (2023b) can also control the location of the subject. We differ significantly from this work in many aspects. First, Anydoor aims at image composition, while our goal is for text-to-image generation. Second, Anydoor can only control the location of the subject, while our work controls the layout of all the objects and regions both in the foreground and background.
@@ 2.2 Grounded Text-to-Image Generation
Given a layout that contains bounding boxes labeled with object categories, grounded text-to-image generation aims to generate the corresponding image, which is the reverse process of object detection. Traditional grounded text-to-image generation such as LostGAN (Sun & Wu, 2019), LAMA (Li et al., 2021) and PLGAN (Wang et al., 2022) are based on generative adversarial networks(GANs). Recently since Stable Diffusion Rombach et al. (2022b), diffusion-based methods (Zheng et al., 2023b; Li et al., 2023; Zhang et al., 2023; Wang et al., 2024b) has tried to add layout control on latent diffusion models, among which LayoutDiffusion (Zheng et al., 2023b) uses patch-based fusion method, GLIGEN (Li et al., 2023) injects grounded embeddings into gated Transformer layers, ControlNet (Zhang et al., 2023) uses copied encoders and zero convolutions, InstanceDiffusion Wang et al. (2024b) allows for multiple formats of location control. However, these methods can only achieve generation based on text prompts, while object-guided generation and identity preservation cannot be achieved. Our framework is an object-centric method, which achieves joint layout guidance of reference object and text, providing flexible and precise layout control for both reference objects and text-guided background generation.
3 Preliminary
Our model is based on Stable Diffusion v1.4, a Latent Diffusion model (LDM) that applies the diffusion process in a latent space. For image generation, it takes the form of a neural network that learns to denoise images that have been augmented with Gaussian noise. Specifically, an input image $x$ is encoded into the latent space using a pretrained autoencoder $z=\mathcal{E}(x),\hat{x}=\mathcal{D}(z)$ (with an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$). Then the denoising process is achieved by training a neural network $\epsilon_{\theta}\left(z_{t},t,v\right)$ which predicts the added noise following:
$\min_{\theta}E_{z_{0},\epsilon\sim\mathcal{N}(0,I),t\sim\mathrm{U}(1,T)}\left\|\epsilon-\varepsilon_{\theta}\left(z_{t},t,v\right)\right\|_{2}^{2}.$ (1)
in which $v$ is the embedding generated from the conditioning variable and $z_{t}$ is the latent noise at timestamp $t$.
In Stable Diffusion Rombach et al. (2022b), the Diffusion UNet is a stack of transformer blocks, which is composed of self-attention and cross attention layers. Specifically, in the cross-attention layer, the features $f$ are updated following:
@@ 3 Preliminary
$Q=\phi_{Q}(f),K=\phi_{K}\left(f_{t}\right),V=\phi_{V}\left(f_{t}\right)$ (3)
@@
where $f_{t}$ is the text embedding, $\phi_{Q}$, $\phi_{K}$, $\phi_{V}$ are the linear layers to project the features into queries, keys and values respectively.
Figure 2: An overview of our proposed framework. It is divided into two steps: (1) Feature extraction. We use CLIP encoder and DINOv2 encoder to extract text and image embeddings respectively, and use our proposed grounding module to extract the grounding tokens. (2) Foreground-background cross-attention control in each transformer block of U-Net. During training, we use dataset with single reference object. In inference stage, the pipeline allows feature injection of multiple reference objects through copied masked cross-attention layers. We show the grounding module in Fig. 3 and masked cross-attention in Fig. 4.
4 Our Approach
Given background-free [^1] Background-free object images can be easily obtained by segmentation methods such as SAM Kirillov et al. (2023), SAM2 Ravi et al. (2024) reference object images $x\in\mathbb{R}^{h\times w\times 3}$ of a subject with the subject’s target spatial locations $L_{i}$, background text entities $T_{j}$ with their target locations $L_{j}$, and the overall prompt $T$, we aim to generate a customized image $\hat{x}$, where both the reference objects can be seamlessly placed inside the desired bounding box with natural poses and accurate identity and the background objects generated from text-box pairs are positioned at the right location. The reference object should be generated harmoniously with the text-generated background. The model can be represented as:
$\hat{x}=\operatorname{GroundingBooth}(x,T,L_{I},L_{T}).$ (4)
where $L_{I}$ represents the layout of the foreground reference objects and $L_{T}$ represents the layout of the background text entities. The pipeline of our proposed GroundingBooth model is shown in Fig. 2. Our work is the first attempt that introduces precise grounding in the customized image synthesis task, which jointly controls the size and location for both the image-driven foreground objects and text-driven background, adaptively harmonizing the poses of the reference objects as well as faithfully preserving their identity. In this section, we first introduce our feature extraction pipeline in sec 4.1, then introduce the foreground-background masked cross-attention control in sec 4.2. Finally we introduce the training and inference pipeline in sec 4.3 and sec 4.4 respectively.
Figure 3: Grounding Module of our proposed framework. Our grounding module takes both the prompt-layout pairs and reference object-layout pairs as input. For the foreground reference object, both CLIP text token and the DINOv2 image class token are utilized.
4.1 Feature Extraction
To maintain accurate identity of the reference object and achieve precise text-driven background generation, we extract both the text tokens from the caption and the image tokens from the references. To control the layout of the foreground and background objects, we design a grounding module, which jointly ground text and image features through position encoding.
Text and Reference Subject Feature Extraction
For the text features, we use CLIP text encoder to extract the text tokens. For the reference subjects, we adopt DINOv2 Oquab et al. (2023), the state-of-the-art self-supervised image feature extractor, and reshape the feature dimension to 768 through a linear projection layer. Specifically, for DINOv2, the extracted 257 ID tokens are composed of a global image class token and 256 local patch tokens. The image class token is extracted separately and paired with the position embedding of the reference object, which is then passed to the grounding module. Here we only extract the image class token since we need to balance the number of image tokens and text tokens in the grounding module.
Grounding Module
Given the entities, which include the foreground reference objects and background text entities, our grounding module aims to jointly ground the foreground and background through positional encoding, which is shown in Fig. 3. We extract grounding information based on the joint guidance of the tagged text-box pair and the object-layout pair. Specifically, this process contains two branches: 1) In the background branch, the bounding boxes of the background objects $L_{T}$ are passed through a Fourier encoder and obtain the text position embeddings; which are then concatenated with the Fourier embeddings in the feature space to get the grounded text embeddings. 2) In the foreground branch, the bounding boxes of the reference object $L_{I}$ (in the target image) are also passed through a Fourier encoder to extract the position embeddings, which are then concatenated with the reference image embeddings to get the grounded reference image embeddings. At the end of the following two branches, the grounded text embeddings and reference image embeddings are reshaped back into the original feature dimension through linear layers and finally concatenated in the token space to get the final grounding embeddings. For each text entity $T$ and reference object entity $I$, grounding manipulation is represented as:
$h^{(T,I)}=\left[\textit{MLP}\left(f_{\text{text }}(T),\textit{Fourier}(L_{T})\right),\textit{MLP}\left(f_{\text{obj }}(I),\textit{Fourier}(L_{I})\right)\right]$ (5)
where Fourier represents the Fourier embedding Tancik et al. (2020), $MLP(.,.)$ is a multi-layer perceptron, [.,.][.,.][ . , . ] is concatenation operation, $L$ is the location information which can be represented as $\bm{L}=\left[x_{\min},y_{\min},x_{\max},y_{\max}\right]$, and $h^{(T,I)}$ is the grounding token. The generated grounding token $\bm{h}^{(T,I)}$ contains the location features of both the reference object and the text prompt. It is then injected into the U-Net layers of the diffusion models. Specifically, inspired by GLIGEN Li et al. (2023), we inject the grounding token through a gated self-attention layer, which is located between the self-attention layer and cross-attention layer in each Transformer block of the U-Net, represented as:
$\bm{v}=\bm{v}+\tanh(\gamma)\cdot\left(\operatorname{SelfAttn}\left(\left[\bm{v},\bm{h}^{(T,I)}\right]\right)\right)$ (6)
where $\gamma$ is a learnable scalar initialized as 00, $\bm{h}^{(T,I)}$ is the grounding token and $v$ is the output of the self-attention layer. During training, the model adaptively learns to adjust the weight $\gamma$ of the grounding module, which ensures stable training and balances the weight between the grounding token and the visual features.
Figure 4: Pipeline of our proposed masked cross-attention. Q, K, and V are image query, key, and value respectively, and A is the affinity matrix.
4.2 Foreground-Background Cross-attention Control
Previous text-to-image generation methods usually directly concatenate the text and image tokens, which leads to two drawbacks. First, the reference foreground object and the text-driven generated background are sometimes blended. Second, for the circumstances that bounding boxes belong to the same class, the model cannot distinguish whether a bounding box belongs to a reference object entity or text entity, which leads to the misplacement of the reference object. To solve these issues, we propose a novel masked-cross attention module to divide the foreground and background while maintaining the independence of generating each foreground object, which jointly grounds each foreground image embedding and background text embedding, shown in Fig. 4.
@@ 4.2 Foreground-Background Cross-attention Control
Specifically, in this layer, both the DINOv2 image tokens and the layout of the reference object are taken as input. We first compute the affinity matrix $A$ through $A=Q\cdot K$ and get $A\in\mathbb{R}^{hw\times hw}$, where $h\times w$ indicates the resolution of the feature map in the attention layer. As we have the object layout $L_{obj}$, it is possible to restrict the injection of image tokens only in the region of the target bounding box. Therefore, we reshape the layout $L_{obj}$ to $h\times w$ and generate the cross attention mask, which is represented as:
$M_{Layout}=\left\{\begin{array}[]{ll}M,&L_{obj}=1,\\ -\infty,&L_{obj}=0\end{array}\right.$ (7)
@@
where $M_{Layout}$ represents the rectified attention score maps, $L_{obj}$ represents the layout region of the reference object in the feature map.
@@
The mask contains the location information for restricting the reference object placement and avoiding information leakage. After acquiring the mask, we apply dot product operation between the feature maps and the layout to constrain the object generation and get the mask-rectified affinity matrix $A^{\prime}$ through $A^{\prime}=A\odot M_{Layout}$. Then we multiply the masked affine matrix $A^{\prime}$ with $V$, and apply mask-guided rectification on the final generated image again to ensure that all the information interactions are restricted in the mask region. The whole masked cross-attention manipulation is represented as:
$O_{obj}=\operatorname{softmax}\left(\frac{QK^{T}\odot M_{Layout}}{\sqrt{d}}\right)V\odot M_{Layout},$ (8)
@@
where $\sqrt{d}$ is the scaling factor that is set as the dimension of the queries and keys, and $\odot$ represents Hadamard product. Through masked cross-attention control, the injection of each reference object feature is restricted into the corresponding bounding box area. This ensures both the independence between foreground and background generation and the independence in the generation of different foreground objects, preventing information leakage and ensuring precise layout-guided reference object generation.

### v2

@@ Abstract
Recent studies in text-to-image customization show great success in generating personalized object variants given several images of a subject. While existing methods focus more on preserving the identity of the subject, they often fall short of controlling the spatial relationship between objects. In this work, we introduce GroundingBooth, a framework that achieves zero-shot instance-level spatial grounding on both foreground subjects and background objects in the text-to-image customization task. Our proposed text-image grounding module and masked cross-attention layer allow us to generate personalized images with both accurate layout alignment and identity preservation while maintaining text-image coherence. With such layout control, our model inherently enables the customization of multiple subjects at once. Our model is evaluated on both layout-guided image synthesis and reference-based customization tasks, showing strong results compared to existing methods. Our work is the first work to achieve a joint grounding on both subject-driven foreground generation and text-driven background generation. The project page is available at https://groundingbooth.github.io..
Figure 1: We propose GroundingBooth, a framework for grounded text-to-image customization. GroundingBooth supports: (a) grounded single-subject customization, and (b) joint grounded customization for multi-subjects and text entities. In general, it achieves a joint grounding on the generation of both the subject-driven foreground and the text-driven background, while preserving the identity of subjects and text-image alignment.
@@ 1 Introduction
Text-to-image customization, also known as subject-driven image generation or personalized text-to-image synthesis, is a task that requires a model to generate diverse variants of a subject given a set of images of the target subject. Text-to-image customization has achieved significant progress during the past few years, allowing for more advanced image manipulation. For example, the test-time-finetuning based methods like Dreambooth (Ruiz et al., 2023), Textual Inversion (Gal et al., 2022), and Custom Diffusion (Kumari et al., 2023) use a few images of the same object to finetune a pretrained diffusion model and generate variants of the object from input prompts. The encoder-based methods such as ELITE (Wei et al., 2023) and InstantBooth (Shi et al., 2023) eliminate test-time-finetuning by learning a generalizable image encoder and attention modules. Despite their success, existing personalization methods mainly focus on generating identity-preserved images from the input prompt and fail to accurately describe the spatial relationship of objects and backgrounds. In real-world scenarios of image customization, it is a crucial user need to achieve fine-grained and accurate layout control on each of the generated objects for more flexible image manipulation.
@@
To tackle this issue, we propose to deal with a more challenging task, grounded text-to-image customization, which extends the existing text-to-image customization task by enabling grounding controllability over both the foreground subjects and background objects. Specifically, under this new setting, the inputs usually include a prompt, images of subjects, and optional bounding boxes of the subjects and background text entities. The model aims to generate text-aligned background objects and identity-preserved foreground subjects, while the spatial location of all the grounded objects and subjects are exactly aligned with the input bounding boxes. It is non-trivial to achieve all these effects simultaneously, as we are indeed handling multiple tasks together and it is challenging for the model to be generalizable to all sub-tasks.
@@
There are a few related works to handle our new task, which, however, show significant limitations. On the one hand, although existing grounded text-to-image diffusion models such as LayoutDiffusion (Zheng et al., 2023) and GLIGEN (Li et al., 2023) have made attempts at spatial controllability, they cannot achieve identity preservation of the subjects. On the other hand, subject-driven image generation methods mainly focus on the identity preservation of the reference objects, while limited attempts have been made on layout control of either subjects or background objects. There is another line of related works (Chen et al., 2023b; Song et al., 2022; 2024) that achieve customized image composition. They can control the location of the input subject under the image composition setting but are neither able to achieve text-to-image synthesis nor control the spatial location of the background contents.
@@
To fully address our new task, we propose GroundingBooth, a general framework for grounded text-to-image customization. Specifically, based on a pretrained text-to-image model, we build a new joint text-image grounding module that encourages both the foreground subjects and background objects to accurately follow the locations indicated by the input bounding boxes. To further enhance the identity preservation of the subjects, we propose a masked cross-attention layer in the transformer blocks of the diffusion U-Net, which helps to disentangle the subject-driven foreground generation and text-driven background generation in each block, effectively preventing the false blending of multiple visual concepts in the same location and enforcing the generation of clear subjects. As shown in Fig. 1, with such dedicated designs of the model structures, our framework not only achieves grounded text-to-image customization with a single subject (Fig. 1 (a)), but also supports multi-subject customization (Fig. 1 (b)), where users can input multiple subjects along with their bounding boxes, and our model can generate each subject in the exact target region with identity preservation and scene harmonization. Meanwhile, our model also allows for the grounding of multiple background objects (Fig. 1 (b)).
@@
  • We propose a general framework, GroundingBooth, that achieves grounded text-to-image customization. Specifically, our model achieves joint layout grounding of both the foreground subjects and the text-guided background, while maintaining accurate identity of the subjects. Furthermore, our model enables the customization of multiple subjects.
@@ 2.1 Subject-driven Image Generation
Subject-driven image generation, also known as personalized text-to-image generation or image customization, aims to generate target images based on customized objects and a text prompt that describes the target context (Chen et al., 2023a; Pan et al., 2024; Xiao et al., 2023; Wang et al., 2024a). In this task, the specific identity of the input reference images is defined as a subject or a concept. So far, existing image customization works can be categorized into three major types. The first type is test-time-finetuning methods (Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023). These methods first finetune a pretrained diffusion model on a few subject images so that the model is adapted to a new identifier token representing the new concept. Then they generate new images from prompts containing the identifier. Such test-time finetuning is computationally intensive. The second type of method is encoder-based customization methods (Arar et al., 2023; Wei et al., 2023; Shi et al., 2023; Zhang et al., 2024), which eliminates the test-time finetuning by learning a generalizable diffusion model that can adapt to new subjects on a large-scale training data. The generalizable model usually contains image encoders that map the subject images into dense tokens and attention modules that integrate vision tokens with text tokens. These methods can achieve much faster image customization during inference, while rely heavily on large-scale multi-view training data, and identity preservation may not be perfect in out-of-distribution cases. The third type of methods (Roich et al., 2021; Gal et al., 2023) is a combination of the first two methods, which first learn an image encoder to extract identity tokens of the input subject and then finetune the model on the subject images for a few steps.
Note that most existing subject-driven image generation methods focus on synthesizing personalized image variants from prompts. They show quite limited performance in controlling the layout of the generated scenes and modeling the spatial relationship between objects. On the contrary, our model performs well not only in generating identity-preserved, text-aligned images but also in controlling the layout of both the subjects and background. A recent related work Anydoor (Chen et al., 2023b) can also control the location of the subject. We differ significantly from this work in many aspects. First, Anydoor aims at image composition, while our goal is for text-to-image generation. Second, Anydoor can only control the location of the subject, while our work controls the layout of all the objects and regions in the foreground and background.
@@ 2.2 Grounded Text-to-Image Generation
Given a layout containing bounding boxes labeled with object categories, grounded text-to-image generation aims to generate the corresponding image, which is the reverse object detection process. Traditional grounded text-to-image generation such as LostGAN (Sun & Wu, 2019), LAMA (Li et al., 2021) and PLGAN (Wang et al., 2022) are based on generative adversarial networks(GANs). Recently, diffusion-based methods (Rombach et al., 2022; Zheng et al., 2023; Li et al., 2023; Zhang et al., 2023; Wang et al., 2024b) have made attempts to add layout control for image generation. For example, LayoutDiffusion (Zheng et al., 2023) uses a patch-based fusion method. GLIGEN (Li et al., 2023) injects grounded embeddings into gated Transformer layers. ControlNet (Zhang et al., 2023) uses copied encoders and zero convolutions. InstanceDiffusion (Wang et al., 2024b) allows for multiple formats of location control. However, all these methods can only perform text-to-image generation, while object-guided generation and identity preservation cannot be achieved. In contrast, our model achieves satisfactory identity preservation on reference-guided image generation.
Figure 2: An overview of our GroundingBooth model. The whole pipeline contains two steps: (1) Feature extraction. We use the CLIP encoder and DINOv2 encoder to extract text and image embeddings, respectively. We use our proposed grounding module to extract the grounding tokens. (2) Foreground-background cross-attention control in each transformer block of U-Net. During training, we use datasets with a single subject as the reference image and only trains a single masked cross-attention layer per transformer block. During inference, our model supports the generation of multiple subjects in their corresponding locations by reusing the same masked cross-attention layer for each subject. This figure shows the inference pipeline of our model. We show the details of the grounding module and masked cross-attention layer in Fig. 3.
3 Our Approach
Our model is built upon Stable Diffusion v1.4 Rombach et al. (2022). Given one or a few background-free [^1] Background-free images refer to images with background removed. They can be easily obtained by segmentation methods such as SAM (Kirillov et al., 2023) or SAM2 (Ravi et al., 2024) reference images $\mathcal{X}=\{x_{1},x_{2},\cdots,x_{m}\}$ where $x_{m}\in\mathbb{R}^{h\times w\times 3}$ with their target bounding box locations $\mathcal{L}_{X}=\{l_{X}^{1},l_{X}^{2},\cdots,l_{X}^{m}\}$, text entities [^2] Here each text entity is referred to a text tag, such as “chair” and “hat”. $\mathcal{T}=\{t_{1},t_{2},\cdots,t_{n}\}$ with their target locations $\mathcal{L}_{T}=\{l_{T}^{1},l_{T}^{2},\cdots,l_{T}^{n}\}$, and the overall prompt $\mathcal{P}$, we aim to generate a customized image $\hat{x}$, where both the reference objects can be seamlessly placed inside the desired bounding box with natural poses and accurate identity and the background objects generated from text-box pairs are positioned at the correct location. Here $l_{X}^{m}$ or $l_{T}^{n}$ refers to the bounding box coordinates of a reference object or a text entity, which can be represented as ${l}=\left[x_{\min},y_{\min},x_{\max},y_{\max}\right]$. The reference object should be generated harmoniously with the background. The customized image $\hat{x}$ can be calculated as:
$\hat{x}=\operatorname{GroundingBooth}(\mathcal{X},\mathcal{T},\mathcal{P},\mathcal{L}_{X},\mathcal{L}_{T}).$ (1)
The pipeline of our proposed GroundingBooth model is shown in Fig. 2. Our work is the first attempt that enables precise spatial grounding in the customized image synthesis task, which jointly controls the size and location for both the reference-guided foreground objects and text-driven background regions. Moreover, our work adaptively harmonizes the poses of the reference objects and faithfully preserves their identity. In this section, we first introduce our feature extraction pipeline in Sec. 3.1, then introduce the foreground-background masked cross-attention control in Sec. 3.2. Finally, we introduce the training and inference pipeline in Sec. 3.3 and Sec. 3.4, respectively.
3.1 Feature Extraction
Feature Extraction of Text and Reference Images
We first extract text tokens from the input prompt using the CLIP text encoder and identity tokens from the reference images using DINOv2 (Oquab et al., 2023). For each reference image, we extract 257 identity tokens which are composed of a global image class token and 256 local patch tokens. We reshape the feature dimension of each image token to 768 through a linear projection layer.
Figure 3: Modules of our proposed framework: (a) Grounding Module: Our grounding module takes both the prompt-layout pairs and reference object-layout pairs as input. For the foreground reference object, both CLIP text token and the DINOv2 image class token are utilized. (b) Masked Cross-Attention: Q, K, and V are visual query, key, and value respectively, and A is the affinity matrix.
Grounding Module
To control the layout of the foreground and background objects, we propose a grounding module, which jointly ground text and image features through positional encoding. Fig. 3(a) shows the overall structure of our grounding module. We extract grounding information based on the joint guidance of the tagged text-box pair and the object-layout pair. Specifically, it contains two branches: 1) In the text entity branch, the bounding boxes of the background objects $\mathcal{L}_{T}$ are passed through a Fourier encoder to obtain the text Fourier embeddings for the text entities, which are then concatenated with the text tokens in the feature space to obtain the grounded text embeddings. 2) In the reference image branch, the bounding boxes of the reference objects $\mathcal{L}_{X}$ (in the target image) are also passed through a Fourier encoder to extract the object Fourier embeddings, which are then concatenated with the reference image tokens to obtain the grounded reference image embeddings. For all training images, we set a max number of boxes and the text entities and drop the rest ones. For the cases where there is no reference image or text entities, we set the input reference object layout to zero and reference object token to zero embeddings, or set the grounded text embeddings to zero embeddings, respectively. At the end of the following two branches, the grounded text embeddings and reference image embeddings are reshaped back into the original feature dimension through linear layers and then concatenated in the token space to form the final grounding tokens. Given the text entities $\mathcal{T}$ and reference images $\mathcal{X}$, the calculation of the grounding tokens is formulated as:
$h^{(\mathcal{T},\mathcal{X})}=\left[\textit{MLP}\left(\psi_{\text{text }}(\mathcal{T}),\textit{Fourier}(\mathcal{L}_{T})\right),\textit{MLP}\left(\psi_{\text{obj }}(\mathcal{X}),\textit{Fourier}(\mathcal{L}_{X})\right)\right],$ (2)
where Fourier represents the Fourier embedding (Tancik et al., 2020), $MLP(.,.)$ is a multi-layer perceptron, [.,.][.,.][ . , . ] is concatenation operation, and $h^{(\mathcal{T},\mathcal{X})}$ is the grounding token. $\psi_{text}$ and $\psi_{obj}$ denote to the text encoder and image encoder, respectively. The generated grounding token ${h}^{(\mathcal{T},\mathcal{X})}$ contains the location features of both the reference objects and the text entities. It is then injected into the U-Net layers of the diffusion models. Specifically, inspired by GLIGEN (Li et al., 2023), we inject the grounding token through a gated self-attention layer located between the self-attention layer and cross-attention layer in each Transformer block of the U-Net, represented as:
${v}={v}+\tanh(\gamma)\cdot\left(\operatorname{SelfAttn}\left(\left[{v},{h}^{(\mathcal{T},\mathcal{X})}\right]\right)\right),$ (3)
where $\gamma$ is a learnable scalar initialized as 00, ${h}^{(\mathcal{T},\mathcal{X})}$ is the grounding token and $v$ is the output of the self-attention layer. During training, the model adaptively learns to adjust the weight $\gamma$ of the grounding module, which ensures stable training and balances the weight between the grounding token and the visual features.
3.2 Foreground-Background Cross-attention Control
Previous text-to-image generation methods usually directly concatenate the text and image tokens in the cross-attention layers, leading to two drawbacks. First, the reference objects and the text-driven background objects can be blended unnaturally. Second, for the circumstances where bounding boxes belong to the same class, the model cannot distinguish whether a bounding box belongs to a reference object or text entity, resulting in the misplacement of the reference object. To solve these issues, we propose a novel masked-cross attention module to separately generate the foreground objects and background contents. Moreover, when there are multiple reference objects, our module can clearly maintain the independence of generating each foreground object. The details of our module is illustrated in Fig. 3(b).
The original cross-attention layer can be formulated as:
@@ 3.2 Foreground-Background Cross-attention Control
$Q=\phi_{Q}(f),K=\phi_{K}\left(f_{c}\right),V=\phi_{V}\left(f_{c}\right),$ (5)
@@
where $\sqrt{d}$ is the scaling factor that is set as the dimension of the queries and keys, $f$ denotes to the vision feature in the layer, $f_{c}$ refers to the embedding of the input condition. $\phi_{Q}$, $\phi_{K}$, $\phi_{V}$ are the linear layers to project the features into queries, keys and values, respectively.
@@
In our masked cross-attention layer, both the DINOv2 image tokens and the layout of the reference object are taken as input. The queries $K$ and values $Q$ are calculated from the image tokens. We first compute the affinity matrix $A$ through $A=Q\cdot K$ and get $A\in\mathbb{R}^{hw\times hw}$, where $h\times w$ indicates the resolution of the feature map in the attention layer. As we have the object layout $l_{obj}$, it is straightforward to restrict the injection of image tokens only inside the region of the target bounding box. Therefore, we reshape the layout $l_{obj}$ to $h\times w$ and generate the cross attention mask, which is formulated as:
$M_{Layout[i,j]}=\left\{\begin{array}[]{ll}0,&[i,j]\in l_{obj},\\ -\infty,&[i,j]\notin l_{obj}\end{array}\right.,$ (6)
@@
where $M_{Layout[i,j]}$ represents the pixel of position $[i,j]$ in rectified attention score maps, $l_{obj}$ represents the layout region of the reference object in the feature map.
@@
The mask contains the location information for restricting the reference object placement and avoiding information leakage. After acquiring the mask, we apply dot product operation between the feature maps and the layout to constrain the object generation and obtain the mask-rectified affinity matrix $A^{\prime}$ through $A^{\prime}=A+M_{Layout}$. Then we multiply the masked affine matrix $A^{\prime}$ with $V$ to obtain the layout-guided masked cross-attention output $f_{obj}$. The whole masked cross-attention module is formulated as:
$f_{obj}=\operatorname{softmax}\left(\frac{QK^{T}+M_{Layout}}{\sqrt{d}}\right)V.$ (7)
@@
For the scenarios where there is a lack of reference objects, $M_{Layout}$ is set to all 00, the masked-cross attention degrade into normal cross attention. Through masked cross-attention control, the injection of each reference object feature is restricted to be within the corresponding bounding box area. This ensures not only the independence between the generation of foreground and background, but also the independence of multiple reference objects. Our module prevents information leakage and ensures an accurate layout-guided subject generation.